OpenStudy (anonymous):

for the functions f(x)= 2x/(x-3), compare the slopes of the tangents i) at the points where x=3.5 and x=20 how do i do this?

OpenStudy (anonymous):

Can you find the first derivative of f(x)

OpenStudy (anonymous):

for x=3.5, it is 14 and for x=20 is 2.36

OpenStudy (anonymous):

That's the slope of tangents at that points. (If you did find f'(x) correctly :))

OpenStudy (anonymous):

the answer says -24 for 3.5 tho :S

OpenStudy (anonymous):

Tho? :? \[f'(x)=\frac{2(x-3)-2x}{(x-3)^2}\\f'(3.5)=-24 \]

OpenStudy (anonymous):

though*.. I dont get it.. what did you do?

OpenStudy (anonymous):

I thought it was some new math symbol. If f(x)=u/v then f'(x)=(u'v-uv')/v^2 Here, u=2x u'=2 v=x-3 v'=1

OpenStudy (anonymous):

@ sabika13 did you get it? any question here?

OpenStudy (anonymous):

use the quotient rule here

OpenStudy (anonymous):

d(u/v)dx=(v du/dx - u dv/dx)/v^2

OpenStudy (anonymous):

can follow from here?

OpenStudy (anonymous):

for y= 2x/(x-3), let u =2x and v=(x-3) now do the quotient rule

OpenStudy (anonymous):

@ sabika13 are you following now here?

OpenStudy (anonymous):

ivenever heard of the quotient rule.. or imforgetting..

OpenStudy (anonymous):

what does d and v stand for :S

OpenStudy (anonymous):

oh i see, are you in algebra class now?

OpenStudy (anonymous):

advanced functions

OpenStudy (anonymous):

yes, do you know limits?

OpenStudy (anonymous):

limits of a function?

OpenStudy (anonymous):

ok advance functions, im sure you already done limits of a function

OpenStudy (anonymous):

im not sure if idid.. or what that is

OpenStudy (anonymous):

\[\lim _{x->0}\frac{ f(x+ \Delta x )-f(x) }{ \Delta x }\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

do that limit as x approach to zero and you will arrived at f'(x)=-6/(x-3)^2 then you plug in x=3.5 and x=20 that is f'(3.5)=-24 and f'(20)=-0.021 and for f(3.5)=14 and f(20)= 3.36

OpenStudy (anonymous):

YW, good luck now